The Food Marketing Institute shows that 17% of households spend more than $100 per week on groceries. Assume the population proportion is p = 0.17 and a sample of 900 households will be selected from the population. Use z-table. 


a. Calculate σ(p̅), the standard error of the proportion of households spending more than $100 per week on groceries to 4 decimals 

b. What is the probability that the sample proportion will be within +/- 0.02 of the population proportion (to 4 decimals)? 

c. What is the probability that the sample proportion will be within +/- 0.02 of the population proportion for a sample of 1,400 households (to 4 decimals)?
Answer #1

given : \(p=0.17, n=900\)

a. \(\sigma_{\hat{p}}=\sqrt{\frac{p *(1-p)}{n}}=\sqrt{\frac{0.17 *(1-0.17)}{900}}=0.0125\)

b. formula : \(z=\frac{\hat{p}-p}{\sqrt{\frac{p^{*(1-p)}{n}}{}}}\)

\(P(\) within \(\pm 0.02) \Rightarrow P(0.15 \(\Rightarrow P\left(\frac{0.15-0.17}{\sqrt{\frac{0.17 *(1-0.17)}{900}}}<\frac{\hat{p}-p}{\sqrt{\frac{p^{*(1-p)}}{n}}}<\frac{0.19-0.17}{\left.\sqrt{\frac{0.17 *(1-0.17)}{900}}\right)}\right.\) \(\Rightarrow P(-1.60 \(\Rightarrow P(z<1.60)-P(z<-1.60)\) \(\Rightarrow 0.9452-0.0548(\) from \(\mathrm{z}-\) table \()\) \(\Rightarrow 0.8904\)

c. formula : \(z=\frac{\hat{p}-p}{\sqrt{\frac{p^{*(1-p)}{n}}{}}}\) \(P(\) within \(\pm 0.02) \Rightarrow P(0.15<\hat{p}<0.19)\) \(\Rightarrow P\left(\frac{0.15-0.17}{\sqrt{\frac{0.17 *(1-0.17)}{1400}}}<\frac{\hat{p}-p}{\sqrt{\frac{p^{*(1-p)}}{n}}}<\frac{0.19-0.17}{\sqrt{\frac{0.17 *(1-0.17)}{1400}}}\right)\) \(\Rightarrow P(-1.99 \(\Rightarrow P(z<1.99)-P(z<-1.99)\) \(\Rightarrow 0.9767-0.0233(\) from \(\mathrm{z}-\) table \()\) \(\Rightarrow 0.9534\)